# Sublinear Data Structures for Nearest Neighbor in Ultra High Dimensions

### v1

@@ D.1 Lower Bound for Approximate Nearest Neighbor
Proof.
The lower bound on the sketch size follows directly from Theorem 2 of [AIP06].
Theorem D.2 (Theorem 2 of [AIP06]).
@@
Assume for contradiction that we have a data structure of size $n^{0.99}$. Then Bob can send this sketch to Alice and let Alice decide the $(1+\epsilon)$-approximate nearest neighbor problem, which contradicts Theorem D.2.
Next, we show a lower bound $\Omega(n)$ of the number of coordinate queries. Consider a set of points $\{c_{i}\}\subset\mathbb{R}^{n}$, where $c_{i}=e_{i}$ is the unit vector. Let the query to be co-located with one of the points, so that the approximate nearest neighbor has to be the exact nearest neighbor. To find out which coordinate of the query is non-zero, a deterministic data structure needs to look at all $n$ coordinates. One can make the same claim about randomized data structures by applying the Yao’s Principle. ∎
D.2 Lower Bound for Exact Nearest Neighbor
@@
Proof.
The proof is a slight modification of Theorem B.1 of [IW18] (with a change in notation which we explain below). First, note that, if Bob can compute $\textsf{cost}(C,X)$ exactly, then it must be the case that the map $\sigma$ that maps each point $x\in X$ to cluster numbered $\sigma(x)$ correctly computes the nearest cluster for each $x$. Indeed, if this is not the case, some $x\in X$ is mapped to a cluster further away from its nearest one. In this case, the additive error incurred cannot be negated by other points in $X$ as none of those points cannot be mapped to a cluster which is closer than its nearest one.
We assume $d=d^{\prime}+1+\log(n)$. Setting $\epsilon=\sqrt{2/d^{\prime}}$ and fixing $\ell=1/\epsilon^{2}$ in Theorem B.1 of [IW18], we have dataset that consists of $2n$ many points, $x_{1},\cdots,x_{n}$ and $z_{1},\cdots,z_{n}$ each of dimension $d$ such that:
@@
Similar to [IW18], we can see that the number of bits required to represent $x_{1},\cdots,x_{n}$ is at least $\log({d^{\prime}\choose d^{\prime}/2}^{n})=d^{\prime}n=\Omega(dn)$ which is our desired lower bound. Also, we note that for a query point $y_{ij}$ whose first $d^{\prime}+1$ coordinates are 0 except the $j$-th coordinate which is set to 1, and the rest of the coordinates are similar to $x_{i}$ or $z_{i}$, the following properties hold:
@@
An exact nearest neighbor query, which can be obtained from $\sigma$ will reveal the value of $x_{i}(j)$ for any $i\in[k]$ and $j\in[d^{\prime}]$. This proves the $\Omega(dn)$ lower bound. ∎

### v2

@@ G.1 Lower Bound for Approximate Nearest Neighbor
Proof G.2.
The lower bound on the sketch size follows directly from Theorem 2 of [4].
Theorem G.3 (Theorem 2 of [4]).
@@
Assume for contradiction that we have a data structure of size $n^{0.99}$. Then Bob can send this sketch to Alice and let Alice decide the $(1+\epsilon)$-approximate nearest neighbor problem, which contradicts Theorem G.3.
Next, we show a lower bound $\Omega(n)$ of the number of coordinate queries. Consider a set of points $\{c_{i}\}\subset\mathbb{R}^{n}$, where $c_{i}=e_{i}$ is the unit vector. Let the query to be co-located with one of the points, so that the approximate nearest neighbor has to be the exact nearest neighbor. To find out which coordinate of the query is non-zero, a deterministic data structure needs to look at all $n$ coordinates. One can make the same claim about randomized data structures by applying the Yao’s Principle.
G.2 Lower Bound for Exact Nearest Neighbor
@@
Proof G.5.
The proof is a slight modification of Theorem B.1 of [25] (with a change in notation which we explain below). First, note that, if Bob can compute $\textsf{cost}(C,X)$ exactly, then it must be the case that the map $\sigma$ that maps each point $x\in X$ to cluster numbered $\sigma(x)$ correctly computes the nearest cluster for each $x$. Indeed, if this is not the case, some $x\in X$ is mapped to a cluster further away from its nearest one. In this case, the additive error incurred cannot be negated by other points in $X$ as none of those points cannot be mapped to a cluster which is closer than its nearest one.
We assume $d=d^{\prime}+1+\log(n)$. Setting $\epsilon=\sqrt{2/d^{\prime}}$ and fixing $\ell=1/\epsilon^{2}$ in Theorem B.1 of [25], we have dataset that consists of $2n$ many points, $x_{1},\cdots,x_{n}$ and $z_{1},\cdots,z_{n}$ each of dimension $d$ such that:
@@
Similar to [25], we can see that the number of bits required to represent $x_{1},\cdots,x_{n}$ is at least $\log(\binom{d^{\prime}}{d^{\prime}/2}^{n})=\Omega(dn)$ which is our desired lower bound. Also, we note that for a query point $y_{ij}$ whose first $d^{\prime}+1$ coordinates are 0 except the $j$-th coordinate which is set to 1, and the rest of the coordinates are similar to $x_{i}$ or $z_{i}$, the following properties hold:
@@
An exact nearest neighbor query, which can be obtained from $\sigma$ will reveal the value of $x_{i}(j)$ for any $i\in[k]$ and $j\in[d^{\prime}]$. This proves the $\Omega(dn)$ lower bound.